# Revela: Dense Retriever Learning via Language Modeling

### v1

@@ 4.1 Experimental Setups
For both retrievers and LMs, we apply LoRA with a rank of 256. Training uses the temperature $\tau$ $1\mathrm{e}{-4}$, the small constant $\epsilon$ $1\mathrm{e}{-6}$, a learning rate of $1\mathrm{e}{-4}$, and 100 warmup steps, following the WarmupDecayLR schedule. We train on 4× A100 80GB GPUs with a gradient accumulation step size of 8. Passages are truncated to 160 tokens, and bf16 mixed precision is enabled. We finetune the models for one epoch, therefore, there are 10,000 steps on Wikipedia and around 11,000 steps on code-related texts.
Specifically, when calculating the similarity between chunks derived from Wikipedia, we follow REPLUG—using only the first half of the chunk to compute similarity with other chunks. This design is motivated by the inferential semantics inherent in natural language retrieval tasks. In contrast, for code retrieval tasks, we retain the full chunk when computing similarity, as these tasks rely more heavily on precise semantic matching compared to natural language.
Table 1: Performance of self-supervised retriever models on BEIR datasets (NDCG@10, %). Bold marks the best score per dataset for unsupervised methods. RepLLaMA (ma2024fine, ), a 1B supervised model trained on MsMarco, is reported for reference only and is not considered when highlighting the best.
Dataset BM25 Contriever RetroMAE LaPraDor REPLUG Revela REPLUG Revela REPLUG Revela Revela+ BM25 RepLLaMA
Model Size 0.1B 0.1B 0.1B 0.1B 0.1B 0.5B 0.5B 1B 1B 1B 1B
ArguAna 48.7 44.3 32.3 44.6 28.1 39.4 34.7 41.4 39.5 45.1 42.2 50.6
C-FEVER 13.6 7.2 5.9 12.2 14.4 15.6 15.3 13.5 14.1 15.6 26.5 27.7
DBPedia 29.9 27.0 4.8 25.0 13.3 18.3 18.9 21.4 19.5 27.6 42.4 42.4
FEVER 48.1 27.2 4.7 33.6 35.1 35.1 47.3 50.9 48.3 60.4 65.7 78.9
FiQA 25.1 12.4 2.8 19.8 14.4 19.1 19.6 26.6 22.0 29.6 41.0 40.1
HotpotQA 56.9 41.0 20.9 30.4 26.7 38.5 36.4 50.6 42.4 56.0 68.8 66.5
NFCorpus 32.1 27.1 6.2 30.4 20.8 23.5 25.8 26.5 26.7 27.2 30.6 35.6
NQ 28.5 18.1 3.7 18.0 18.3 21.2 26.5 29.8 27.9 33.9 44.4 57.8
Quora 80.4 83.4 56.8 78.7 72.3 80.9 76.4 83.1 82.4 83.5 86.7 84.4
SCIDOCS 15.8 10.9 4.1 13.4 9.6 12.1 13.5 14.8 14.5 16.3 27.0 17.8
SciFact 68.7 59.1 30.9 49.9 55.9 57.9 63.9 65.2 67.1 71.8 75.9 74.3
TREC-C 62.2 18.2 15.7 22.9 48.2 60.2 44.5 58.4 39.6 59.8 26.7 83.1
Touche 33.1 7.2 0.4 8.9 11.6 14.4 14.5 21.0 13.5 23.3 47.8 32.8
Mean 41.8 29.5 14.6 29.8 28.4 33.6 33.6 38.7 35.2 42.3 48.1 53.2
Table 2: Performance on code-related retrieval (NDCG@10, %). Gray indicates supervised models trained on code-specific data. Bold marks the best score among all non-proprietary models.
Dataset BM25 Contriever GCBERT UniXcoder REPLUG Revela REPLUG Revela REPLUG Revela Ada-2 Voyage-2
Model Size 0.1B 0.1B 0.1B 0.1B 0.1B 0.5B 0.5B 1B 1B
Apps 4.8 5.2 0.1 1.4 5.2 6.3 9.5 19.0 13.9 16.8 8.7 26.5
CosQA 15.6 13.2 3.6 25.1 16.5 24.2 17.7 27.6 20.1 27.7 28.9 29.8
ST2SQL 25.0 38.6 10.4 50.5 45.0 44.7 53.4 53.3 53.9 53.4 58.3 69.3
SN 40.9 33.2 3.0 60.2 35.0 43.2 42.8 52.6 50.8 60.1 74.2 81.8
SNCC 54.0 46.8 18.3 58.4 52.7 58.9 59.5 63.6 62.8 67.0 53.3 72.8
TransC 47.8 42.5 12.8 41.8 57.0 60.6 53.5 71.3 61.5 74.4 26.0 27.3
TransDL 34.4 33.6 31.7 31.0 34.0 32.9 31.2 34.1 33.3 33.7 17.7 28.7
SOQA 70.2 59.9 20.3 44.7 62.9 68.3 72.4 81.0 76.3 83.8 74.2 81.8
F-ST 68.1 44.5 6.9 36.0 51.1 65.0 59.4 73.8 66.0 74.5 47.1 65.4
F-MT 59.2 38.7 13.1 24.2 31.4 43.1 37.9 52.3 42.9 54.6 17.7 28.7
Mean 42.0 35.6 12.0 37.3 39.1 44.7 43.7 52.9 48.2 54.6 45.6 56.3
@@ 4.2 Experimental Results
Table 1 and Table 2 report the performance of Revela and baseline methods on BEIR and CoIR, highlighting Revela’s superior effectiveness. Specifically, with a comparable number of parameters (around 0.1B), Revela achieves an absolute gain of 4.1% on BEIR and 9.1% on CoIR over Contriever, one of the most popular self-supervised retriever. Moreover, Revela’s performance continues to improve as the retriever size increases to 0.5B and 1B.
Revela consistently outperforms REPLUG across models of varying sizes. On the BEIR benchmark, Revela -1B achieves an average score of 42.3 across 13 datasets, compared to 35.2 for REPLUG-1B. Similarly, on the CoIR benchmark, Revela 1B achieves 54.6, surpassing REPLUG’s 48.2. Given the large evaluation sets (33K queries for BEIR and 162K for CoIR), the improvements are statistically significant, with a z-test yielding a p-value below $10^{-9}$.
Combining the retrieval results of BM25 and Revela using Reciprocal Rank Fusion (RRF) (cormack2009reciprocal, ) effectively narrows the gap with the supervised model RepLLaMA, improving NDCG@10 by 5.8 points. On out-of-domain tasks such as SciDocs, SciFact, and Quora, this combination even surpasses RepLLaMA, highlighting the complementary strengths of BM25 and Revela. For the details about RRF implementation, please refer to Appendix A.2
In Table 2, Revela is shown to beat some supervised models across different sizes on CoIR and some proprietary models. Although Revela does not yet match the performance of Voyage-Code-002—the strongest model benchmarked on CoIR—we anticipate that scaling up the retriever backbone will further narrow the gap, and potentially surpass the closed-source models.
@@ 5 Analysis
Beyond Revela’s effectiveness shown earlier, we will explore its scalability from two aspects: LM size (Sec. 5.1) and training batch configuration (Sec. 5.2). We also validate the impact of our specific design on batch construction, batch sizes (Sec. 5.2), and model architecture (Sec. 5.3).
5.1 Larger LMs can help out-of-domain retrieval tasks
As observed in the previous section, retriever performance improves with larger backbone models. In this section, we further investigate how the size of the LM—another key component in training—affects the retriever’s effectiveness. In addition to LLaMA-3.2-1B, we also use SmolLM2-135M and Qwen2.5-0.5B as LMs, each paired with retrievers of three different sizes. All models are trained using the same experimental setup described in Sec. 4, on both training corpora.
As shown in Fig. 3, larger LMs do not show a consistent advantage in retrieval performance on BEIR–the general-domain benchmark. However, a clear positive trend is observed on the CoIR benchmark. We expect that seamlessly incorporating larger LMs will further enhance Revela’s adaptability to specific domains. For detailed results on each dataset, as well as an analysis of trainable parameters (i.e., LoRA rank) as another aspect of training size, please refer to Appendix B.1.
Figure 3: Performance comparison on BEIR and CoIR benchmarks using various combinations of retrievers and LMs. For code retrieval tasks, larger LMs consistently yield greater gains in retriever performance, whereas this trend is not observed for general-domain retrieval.
5.2 Training Dataset Configuration
Figure 4: Average performance (NDCG@10, %) on the BEIR and CoIR benchmarks across different dataset configurations (vanilla, randomly constructed, and mixed training corpus). Best results are highlighted in bold.
BEIR
Model 135M 500M 1B
Revela 33.6 38.7 42.3
Revela-random 3.9 4.8 4.5
Revela-dual 34.4 38.9 42.9
CoIR
Model 135M 500M 1B
Revela 44.7 52.9 54.6
Revela-random 0.2 0.2 4.8
Revela-dual 43.9 52.2 54.4
Figure 5: Average performance (NDCG@10, %) with and without V-normalization across the retrievers in different sizes. The retrievers with V-normalization perform better than those without the mechanism, signifying the significance of this design.
We investigate the impact of dataset construction strategies, focusing on both batch formation and the composition of the training data.
Batch construction by chunking documents matters.
As previously described, the Revela configuration groups continuous chunks from one or more passages into a single batch, based on the intuition that such chunks are not semantically identical but are contextually related. To assess the importance of this design, we take the training data described in Sec. 3, shuffle all chunks across batches, and regroup them into new, randomly constructed batches. We repeat the same training procedure on both the Wikipedia and code-related corpora to obtain the Revela-random models. As shown in Table 5, the substantial performance drop of Revela-random compared to Revela underscores the importance of structured batch construction. For more detailed results in BEIR and CoIR, please refer to Appendix B.2.
@@ 5.2 Training Dataset Configuration
The in-batch attention mechanism in Revela is inspired by the concept of in-batch negatives in supervised contrastive learning xiong2020approximatenearestneighbornegative . As described in Sec. 4.1, the default batch size is set to 16. To analyze the impact of batch size, we construct training data with batch sizes of 8 and 12 using the same batch construction strategy. To ensure a fair comparison, we maintain a consistent total number of training tokens by proportionally increasing the training steps by 100% and 33.3%, respectively. As shown in Fig. 6, Revela achieves the best performance on both BEIR and COIR when trained with a batch size of 16, consistent with the underlying motivation of in-batch negatives. For more detailed results within BEIR and CoIR, please refer to Appendix B.3.
@@
Figure 6: Performance comparison on BEIR and CoIR benchmarks using different batch sizes. For both benchmarks, Revela trained with the batch size of 16 outperforms those trained with 8 and 12.
Revela efficiently learns from mixed-domain composition of the training corpus.
As discussed in Sec. 4.2, when trained separately on Wikipedia and code-related corpora, Revela achieves strong performance on both general and code retrieval benchmarks (BEIR and CoIR, respectively). To explore the potential for scaling up, we examine how combining datasets from mixed domains affects model performance. Specifically, we randomly sample 160,000 batches each from the general and code-related training sets constructed in Sec. 4.1, resulting in a combined set of 320,000 batches. We retrain Revela on the mixed-domain corpus with the same configuration, especially including the number of training steps, and refer to the resulting model as Revela-dual. As shown in Table 5, Revela-dual performs comparably to or even better than the vanilla Revela across different retriever sizes on both BEIR and CoIR, demonstrating the model’s potential for effective generalization across domains.
Figure 7: Comparison of Revela  trained on different recipes.
We further evaluate Revela’s generalization ability by augmenting the Wikipedia data used in Sec. 4.1 with the BEIR corpus. Specifically, we select five domain-specific datasets, including FiQA, NFCorpus, SciFact, SciDocs, and TRECCOVID, and generate a total of 54,312 training batches. These are combined with the existing Wikipedia batches to form a dataset of 320,000 batches. As shown in Fig. 7, models trained on both Wikipedia and BEIR consistently outperform those trained solely on Wikipedia, demonstrating Revela’s effectiveness in adapting to new domains using only the corpus.
5.3 Ablation of V-normalization
In Section 3, we propose a specific design of Revela for in-batch attention computation, which incorporates V-normalization, as defined in Equation 6. The underlying intuition is to prevent the disproportionate amplification of individual patterns within the sequence (e.g., specific tokens or words), thereby encouraging a more comprehensive representation of the overall semantics.
To validate its effectiveness, we conduct an ablation study by jointly training retrievers and LMs without V-normalization. As shown in Figure 5, removing V-normalization leads to a consistent and significant drop in performance across all three Revela variants with different backbone sizes on the BEIR benchmark. These results highlight the critical role of V-normalization in Revela. For a detailed breakdown of per-dataset results, please refer to Appendix B.4.
6 Discussion
Future Directions
@@ 6 Discussion
This work shows how to train a retriever via self-supervised language modeling on raw text. We outline several future directions to suggest follow-up research. (1) Iterative Indexing: Sec.  5.2 emphasizes the importance of semantically related passages in Revela’s batch construction. While Revela uses document chunking, a more general approach would iteratively index documents and group chunks by on-the-fly representations. Though explored in prior work (izacard2023atlas, ; shi-etal-2024-replug, ), the high computational cost of such methods remains a key challenge for future work. (2) Scaling up: We envision several directions to scale up Revela, including expanding data coverage, as discussed in Sec. 5.2, and increasing model sizes, as analyzed in Sec. 5.1. Additionally, incorporating advanced attention mechanisms (yuan2025native, ) may further enhance retriever learning and accelerate the training process. (3) Potential Influence on LM Training: Revela focuses on retriever learning, treating the LM as auxiliary. LM performance is not analyzed, and it remains unclear whether in-batch attention affects it. Batch construction significantly impacts convergence and model quality (wang2024greats, ; hong2024diversified, ). Future work could explore using retrievers to guide batch construction for LM training.
Limitations
We conclude the limitations of the works from the following perspectives. (1) Evaluation: We only evaluate our methods on an English retrieval benchmark (BEIR) and code retrieval (CoIR). More experiments on multilingual retrieval tasks can make the work more comprehensive. (2) Model scales: As mentioned in Sec. 6, the experiments on larger models (both retrievers and LMs) can better verify the generalization of Revela. (3) Dataset construction: As shown in Sec. 5.2, Revela depends on the intrinsic structure within natural language. How to better leverage the structure for retriever learning is an open question, which has not been covered in this work.
@@ 7 Conclusion
Efficiently building information-seeking systems is crucial due to the swiftly evolving world and the wide existence of specific domains. In this work, we introduce Revela, a self-supervised framework that couples dense retrieval with language modeling through a novel in-batch attention mechanism, where a token attends to both local context and other sequences in the batch during NTP. By letting the retriever’s relevance scores weight cross-sequence attention, Revela transforms NTP into a retrieval signal, making use of high-quality and diverse raw text and eliminating the need for manual relevance labels. Our experiments on both general and domain-specific benchmarks demonstrate significant performance gains over existing self-supervised methods, with improvements scaling with retriever size. Further analysis on training data composition and LM size highlights the potential of Revela as a strong and scalable alternative to traditional self-supervised retriever training paradigms, paving the way for more general and efficient retriever learning.
Acknowledgments and Disclosure of Funding

### v2

@@ 4.1 Experimental Setups
For both retrievers and LMs, we apply LoRA with a rank of 256. Training uses the temperature $\tau$ $1\mathrm{e}{-4}$, a learning rate of $1\mathrm{e}{-4}$, and 100 warmup steps, following the WarmupDecayLR schedule. We train on 4× A100 80GB GPUs with a gradient accumulation step size of 8. Passages are truncated to 160 tokens, and bf16 mixed precision is enabled. We finetune the models for one epoch, namely, there are 10,000 steps on Wikipedia (~ 44 hours) and around 11,000 steps on code-related texts (~ 48 hours). During inference, the max token length of queries and documents is 2048. For more details of the experimental setups, please refer to § B.5.
@@ 4.2 Experimental Results
Revela exhibits superior performance on domain-specific retrieval.
Tab. 1 reports the performance of Revela and baseline methods on CoIR. As an unsupervised model trained without query-document pairs, Revela3B surpasses E5-Mistral-7B-Instruct, a much larger supervised model pre-trained and fine-tuned on massive, well-curated text-doc pairs, as well as two proprietary APIs averaged on 10 tasks. Revela also follows scaling laws: its performance consistently improves with model size, while maintaining superiority at every scale over baselines. At 0.1B parameters, Revela outperforms the code-specific supervised model UniXCoder by 11.1 points on NDCG@10. At the 0.5B scale, our model outperforms E5-PT by nearly 10 points, despite E5-PT being pre-trained on 270 million filtered query-document pairs covering Revela’s corpus. Revela0.5B even surpasses the supervised model BGE-M3, despite the latter being pre-trained on extensive text-code pairs. [^4] Please refer to Table 8 in the original paper (Chen et al., 2024). Moreover, at each scale, Revela also surpasses REPLUG, underscoring the effectiveness of the co-training paradigm between retrievers and LMs.
Table 1: Performance on CoIR (NDCG@10, %). Gray indicates supervised models. Bold marks the highest score among non-API models in each row. Columns marked used code-related pairs during pre-training. The results of APIs are collected from Li et al. (2024). Without query-document pairs, Revela3B surpasses larger supervised models and proprietary APIs, averaged across 10 tasks.
Dataset BM25 UniX Revela E5-PT Revela BGE REPLUG Revela REPLUG Revela E5-Mistral Ada-2 Voyage-2
Model Size 0.1B 0.1B 0.3B 0.5B 0.6B 1B 1B 3B 3B 7B
Apps 4.8 1.4 8.2 10.6 20.5 14.7 13.9 19.4 17.7 26.6 23.5 8.7 26.5
CosQA 15.6 25.1 26.2 27.1 27.5 26.4 20.1 30.2 25.2 29.0 33.2 28.9 29.8
ST2SQL 25.0 50.5 45.7 48.9 53.7 46.9 53.9 55.0 56.8 55.9 68.0 58.3 69.3
SN 40.9 60.2 49.9 35.2 57.9 58.3 50.8 64.0 53.7 62.6 67.4 74.2 81.8
SNCC 54.0 58.4 63.4 50.5 68.0 53.7 62.8 70.0 62.8 69.1 64.8 53.3 72.8
TransC 47.8 41.8 70.9 56.3 77.6 62.6 61.5 81.1 62.1 83.2 80.6 26.0 27.3
TransDL 34.4 31.0 34.6 32.2 35.4 30.2 33.3 34.2 34.4 34.5 31.7 17.7 28.7
SOQA 70.2 44.7 69.2 86.9 82.5 80.6 76.3 85.7 78.1 88.3 91.0 74.2 81.8
F-ST 68.1 36.0 63.8 70.4 74.5 69.3 66.0 76.2 71.7 78.8 76.4 47.1 65.4
F-MT 59.2 24.2 51.7 46.2 63.6 47.9 42.9 70.4 49.1 73.0 36.4 17.7 28.7
Mean 42.0 37.3 48.4 46.4 56.1 49.1 48.2 58.6 53.9 60.1 57.3 45.6 56.3
Figure 3: Performance on BRIGHT (Left) and BEIR (Right) (NDCG@10,%). Results for Revelaare shown in opaque bars, while all other models are represented by transparent bars. On BRIGHT, a reasoning-intensive task, Revela performs on par with large supervised models and properties APIs. On BEIR, Revela achieves similar performance with E5-PT with much less data and compute. Please refer to Tab. 6 and Tab. 7 in § B.6 for the per-task results.
Revela demonstrates strong performance on complex retrieval.
Fig. 3 (Left) shows the average NDCG@10 across the 12 BRIGHT subtasks. Trained unsupervised on only 340K raw Wikipedia documents, Revela8B performs on par with the supervised E5-Mistral-7B-Instruct and other proprietary models, and even surpasses cohere-embed-english-v3.0. The scaling trends observed in code retrieval persist here as well: even Revela0.5B outperforms REPLUG3B with much fewer parameters. At a comparable parameter scale, Revela0.5B exceeds E5-PT by 1.1 points (10% relative), a noteworthy result given the potential advantage from E5-PT’s training corpus overlapping with BRIGHT. These findings highlight Revela’s promise for tackling more complex retrieval scenarios.
Revela achieves efficient and robust generalization across tasks.
Fig. 3 (Right) reports the average NDCG@10 across the 13 BEIR tasks, where Revela proves to be remarkably efficient and robust. At the 0.1B scale, Revela outperforms Contriever and LaPraDor by more than 3 absolute points. Moreover, Revela’s consistently surpass REPLUG by a significant margin (3B: 8.9%; 1B: 7.0%), mirroring trends observed on CoIR and BRIGHT. Remarkably, despite requiring roughly 1000× less training data and over 10× fewer compute resources, Revela3B achieves performance on par with E5-PT, underscoring its efficiency and effectiveness. Notably, even when trained solely on a code-related corpus, Revela3B performs comparably on the general-domain BEIR benchmark to both its Wikipedia-trained counterpart and E5-PT, underscoring its powerful generalization capabilities.
@@ 5 Analysis
We conduct several targeted analyses to further investigate Revela. First, to isolate its algorithmic contribution, we compare it with Contriever using an identical LM backbone. This comparison reveals that Revelaachieves superior performance and exhibits stronger domain-specific robustness to training data. Second, we demonstrate that, consistent with traditional contrastive learning, Revelabenefits from larger batch sizes. Finally, we examine the LM’s impact on retriever performance within the co-training framework. Additional studies on mixed-domain training, out-of-domain generalization, the LM’s post-Revela capabilities, and computational efficiency are presented in Appendix App. C.
Controlled experiments under the same LM backbone.
As baseline models may use different sizes and architectures of LMs as retriever backbones, we implement a classical unsupervised retriever learning algorithm, Contriever (Izacard et al., 2022), using the same model (LlaMA-3.2-1B) and the same datasets (Wikipedia and the code-related corpus introduced in § 4.1) as training data. To construct pseudo query-document pairs from raw text, Contriever mainly applies two tricks: Inverse Cloze Task, where a sentence is removed from a passage to form a query against the remainder, and Independent Cropping, where two spans from the same document form a positive pair while spans from different documents serve as negatives. In this way, we generate 500K and 359K query-document pairs, each with 15 negatives, for general and code domains, respectively. The models contrastively trained on them are denoted as Contriever-wiki1B and Contriever-code1B. Please refer to § C.1 for more training details.
Table 2: Revela vs. Contriever Performance.
Model BEIR CoIR AVG
Revela-wiki1B 42.7 53.2 48.0
Contriever-wiki1B 42.4 50.3 46.4
Revela-code1B 39.6 58.6 49.1
Contriever-code1B 32.3 52.1 42.2
As shown in Tab. 2, Revela outperforms Contriever on both BEIR and CoIR, irrespective of whether training is conducted on general or code-specific data. Moreover, the performance disparity becomes more pronounced in out-of-distribution domains, underscoring Revela’s robustness and its strong capacity for cross-domain generalization over contrastive learning.
@@
The in-batch attention mechanism in Revela is inspired by the concept of in-batch negatives in supervised contrastive learning (Xiong et al., 2020). As described in § 4.1, the default batch size is set to 16. To analyze the impact of batch size, we construct training data with batch sizes of 4 and 8 using the same batch construction strategy. As shown in Fig. 4, Revela’s performance scales with batch size, suggesting potential for further gains. For more detailed results within CoIR and BEIR, please refer to § C.2.
@@
Figure 4: Performance comparison on CoIR and BEIR with different batch sizes. For both benchmarks, Revela performance generally scales with batch size.
Larger LMs can help out-of-domain retrieval tasks.
As observed in the previous section, retriever performance improves with larger backbone models. In this section, we further investigate how the size of the LM, the other key component in training, affects the retriever’s effectiveness. In addition to LLaMA-3.2-1B, we also use SmolLM2-135M and Qwen2.5-0.5B as LMs, each paired with retrievers of three different sizes. All models are trained using the same experimental setup described in § 4, on both training corpora.
As shown in Fig. 5, CoIR exhibits a clear positive trend: the largest LM delivers best retrieval performance. By contrast, on BEIR, the general-domain benchmark, larger LMs do not provide a consistent advantage. We therefore expect that seamlessly incorporating larger LMs will further improve Revela’s adaptability to specialized domains. For per-dataset results, see § C.3.
Figure 5: Performance comparison on CoIR and BEIR using various combinations of retrievers and LMs. For code retrieval tasks, larger LMs can yield greater gains in retriever performance.
To further investigate Revela, we conducted extra experiments, with the following key findings:
  • Revela learns efficiently from mixed-domain training corpora. Mixing Wikipedia with the code-related corpora used in § 4 maintains, or even improves, retrieval performance, demonstrating Revela’s strong generalization across diverse domains with only raw texts (See § C.4).
  • Even when trained on an LM-training, out-of-domain corpus, Revela still performs competitively, underscoring its robustness and confirming the observation in Tab. 2. At a similar scale, Revela, trained on out-of-domain data, can still outperform E5-PT on CoIR (See § C.5).
  • The co-trained LM’s capacity is largely preserved, plausibly due to the use of LoRA and retaining the NTP objective while adding only auxiliary in-batch attention (See § C.6).
  • Revela offers a theoretical advantage in efficiency over REPLUG, complementing its empirically demonstrated performance gains (See § C.7).
6 Future Directions
@@
We train retrievers directly from raw text via self-supervised language modeling, sidestepping query-doc pairs, a breakthrough beyond the traditional paradigm. Based on this novel solution, We outline several future directions to suggest follow-up research. (1) Iterative Indexing: While Revela uses document chunking, a more general approach would iteratively index documents and group chunks by on-the-fly representations. Though explored in prior work (Izacard et al., 2023; Shi et al., 2024), the high computational cost of such methods remains a key challenge for future work. (2) Scaling up: We envision several directions to scale up Revela, including increasing retriever size in § 4.2, and increasing LM size in § 5. Additionally, incorporating advanced attention mechanisms (Yuan et al., 2025) may enhance retriever learning and accelerate the training. (3) Multi-modality: Although Revela targets text and code retrieval, this retriever-via-language-modeling paradigm can, in principle, generalize to other LM-modeled modalities, such as images (Jiang et al., 2025).
@@ 7 Conclusion
Efficiently building information-seeking systems is crucial due to the swiftly evolving world and the wide existence of specific domains, where query-document curation is one key bottleneck. In this work, we introduce Revela, a self‑supervised framework that couples dense retrieval with language modeling through a novel in‑batch attention mechanism, where a token attends to local context and other sequences in the batch during NTP. By letting the retriever’s relevance scores weight cross‑sequence attention, Revela transforms NTP into a retrieval signal, making use of raw text and eliminating the need for annotation or synthesis. Our experiments on domain-specific, complex, and general benchmarks demonstrate significant performance gains over existing self-supervised methods, with improvements scaling with retriever size. Further analysis on batch size, LM size, and mixed-data composition highlights Revela as a strong and scalable alternative to traditional self-supervised paradigms, paving the way for more general and efficient retriever learning.
Reproducibility statement
To ensure full reproducibility, we provide comprehensive details on our methods and experiments. The setups for our main results, presented in § 4.2, are described in § 4.1, including hyperparameters, training resources, training time, etc. Further implementation details are located in the appendices, covering our model, Revela  (§ B.5), REPLUG implementation (§ B.3) with its checkpoints (§ B.4), evaluation benchmarks (§ B.1), and the training corpus (§ B.2).
The appendices also contain our extended analyses. These include a detailed comparison with unsupervised contrastive learning methods (§ C.1), a study on mixed-domain composition within the training data (§ C.4), and an outline of data usage for out-of-domain generalization (§ C.5). The code to reproduce our results is included in the supplementary submission.
Acknowledgments